# Safe Multi-Robotic Arm Interaction via 3D Convex Shapes

### v2

@@ 7.1 Simulation studies
We conducted our simulation experiments in MuJoCo [28]. The simulation environment, consisting of four Franka Research 3 robotic arms with intersecting workspaces, is illustrated in Fig. 2. Each robot’s end effector and last three links were enclosed by four ellipsoids to define HOCBFs, resulting in a total of 96 HOCBFs for multi-robot safety. In the simulation, we assumed that HOCBF calculations were performed in real-time, ensuring no control delays.
@@
As shown in Tab. 2, the estimated Hessian contribution preserved the HOCBF safety guarantees, indicating its reliability when analytical computation is infeasible. While the centralized filters succeeded in all experiments, the decentralized filters failed in some cases due to QP infeasibility. Although relaxation in decentralized filtering mitigated this issue in some experiments, it did not fully eliminate it. In all instances where a QP became infeasible, the backup controller successfully halted all robots and prevented collisions. When centralized filters were employed as a backup strategy in these cases, we observed that infeasibility did not arise, allowing the task to continue without interruption. Moreover, although the reduced feasible set in the decentralized filter led to slightly more conservative behavior compared to the centralized filters, task execution remained feasible. The relaxed-decentralized filter helped to alleviate this conservatism.

### v3

@@ 7.1 Simulation studies
We conducted our simulation experiments in MuJoCo todorov2012mujoco . The simulation environment, consisting of four Franka Research 3 robotic arms with intersecting workspaces, is illustrated in Fig. 2. Each robot’s end effector and last three links were enclosed by four ellipsoids to define HOCBFs, resulting in a total of 96 HOCBFs for multi-robot safety. In the simulation, we assumed that HOCBF calculations were performed in real-time, ensuring no control delays.
@@
As shown in Tab. 2, the estimated Hessian contribution preserved the HOCBF safety guarantees, indicating its reliability when analytical computation is infeasible. While the centralized filters succeeded in all experiments, the decentralized filters failed in some cases due to QP infeasibility. Although relaxation in decentralized filtering mitigated this issue in some experiments, it did not fully eliminate it. In all instances where a QP became infeasible, the fail-safe mode successfully halted all robots and prevented collisions. When centralized filters were employed as a backup strategy in these cases, we observed that infeasibility did not arise, allowing the task to continue without interruption. Moreover, although the reduced feasible set in the decentralized filter led to slightly more conservative behavior compared to the centralized filters, task execution remained feasible. The relaxed-decentralized filter helped to alleviate this conservatism.